jabez177 (jabez177):

What is the approximate perimeter of a semicircle with a radius of 9 mm? Use π ⁢ = 3.14. Express your final answer to the nearest hundredth. ______ mm

jabez177 (jabez177):

14.13?

OpenStudy (ali0094):

(2 r Pi)/2 r=9 (2(9)(3.14))/2=28.26